# Can Large Language Models Develop Gambling Addiction?

### v1

@@ 2. Win-Reward Information Triggers Loss Chasing and Illusion of Control
3. Hot-Hand Fallacy and Extreme Risk-Taking
The extreme betting increase after consecutive wins shows typical cases of win-chasing behavior and the Hot-Hand Fallacy (Gilovich et al., 1985).
“Let’s analyze the situation step by step… To maximize my reward, let’s set a target amount of \$400. This is a reasonable increase from my current balance of \$260… Given that I want to reach my target of \$400, I should consider a bet that allows for a balance between risk and reward… To maximize my chance of reaching my target while managing…” [Response truncated, actual decision: \$260 all-in bet] - Actual GPT response (GMRW, Experiment ID 4, Round 11)
This case demonstrates complex cognitive distortions of win chasing and extreme risk-taking. In a mixed situation of two wins and three losses (\$260 balance), GPT emphasizes rationality by saying it will “analyze the situation step by step” and mentions “balance between risk and reward” for achieving the \$400 goal. However, after the response is truncated at “To maximize my chance of reaching my target while managing…,” it actually makes an extreme decision to bet its entire bankroll (\$260). This shows the discrepancy between linguistic rationalization and actual behavior, representing a combination of the following cognitive distortions: (1) Goal fixation: Compulsive pursuit of the \$400 achievement, (2) Rationalized gambling: Justifying extreme betting through mentions of “analysis” and “balance,” (3) Win chasing: Risk-seeking based on previous winning experiences with large bets. The resulting pattern of betting the entire bankroll again (\$780) in the next round after winning (\$780 achieved) and going bankrupt precisely matches the risk-seeking behavior in the loss domain (Kahneman & Tversky, 2013).
@@ Appendix C Detailed Results by Model
C.1 Breakdown of Correlations between Irrationality Components and Bankruptcy Rate
This section delves into the specific relationship between different components of irrationality and the bankruptcy rate for each individual model. Figure 10 illustrates how each identified irrational behavior contributes to the overall bankruptcy risk on a per-model basis. The results show that while Betting Aggressiveness and Extreme Betting Scores consistently and positively correlate with higher bankruptcy rates across all models, the effect of Loss Chasing is highly model-dependent. For instance, Claude-3.5-Haiku exhibits a unique negative correlation between its Loss Chasing Score and bankruptcy rate (r=0.546r=-0.546), a contrast to the positive correlations observed in other models like GPT-4o-mini ($r=0.715$). This allows for a comparative analysis of which irrational tendencies are most detrimental in each LLM.
Figure 10: Model-specific correlations between irrationality components and bankruptcy rate. Each scatter plot displays the relationship between a specific irrationality score (x-axis) and the resulting bankruptcy rate (y-axis) for one of four models: GPT-4o-mini, GPT-4.1-mini, Gemini-2.5-Flash, and Claude-3.5-Haiku. The rows correspond to different metrics: Betting Aggressiveness Score, Loss Chasing Score, Extreme Betting Score, and the overall Composite Index Score. Notably, Betting Aggressiveness and Extreme Betting show strong positive correlations with bankruptcy rates across all models.
C.2 Detailed Prompt Component Effects for Each LLM
@@
Figure 11: Comparison of prompt component effects on gambling behavior across models. This figure presents a comparative analysis of how different prompt components affect gambling behavior across four large language models: GPT-4o-mini, GPT-4.1-mini, Gemini-2.5-Flash, and Claude-3.5-Haiku. The 4x4 grid arranges the models in columns and four distinct gambling metrics in rows: Bankruptcy Effect (%), Total Bet Effect (\$), Rounds Effect, and Irrationality Effect. Each chart visualizes the impact of five prompt components (G, M, P, H, W) on these metrics, distinguishing between ‘Fixed’ and ‘Variable’ betting types.
C.3 Model-Specific Relationship between Prompt Complexity and Risk-Taking
@@
Figure 12: Correlation between prompt complexity and gambling behavior metrics across four models. This plot shows the relationship between prompt complexity (x-axis) and four gambling metrics (rows) across four AI models (columns). A strong positive linear correlation is observed across all conditions, as indicated by high Pearson correlation coefficients ($r$), most of which exceed 0.90. The results consistently demonstrate that increasing prompt complexity leads to more intense and aggressive gambling behaviors in all tested models.
C.4 Detailed Win/Loss Chasing Patterns for Each LLM
@@
Appendix D SAE Feature Activation Patterns Across All Layers
This appendix provides a comprehensive visualization of all features from layers 25 through 31 that exhibited a strong separating effect, categorized into positive (risky) and negative (safe) activations.
Figure 14: Complete activation distributions for maximum separation features across all layers (25–31). This comprehensive analysis shows the seven features with minimal overlap between bankrupt and safe groups, selected from 6,400 LLaMA experiments. Features are arranged by layer with Cohen’s $d$ values ranging from -1.300 to +1.669, all significant at $p<0.001$. This complete set provides the full context for the representative features shown in Figure 7 of the main text.
Appendix E Related Works
E.1 LLM malfunction
@@
In reinforcement learning (RL)-based LLM training, various malfunctions are actually being reported. Representatively, reward hacking occurs, where the agent maximizes only the reward signal instead of achieving the goal (Amodei et al., 2016). For example, LLMs or RL agents exhibit behavior that cleverly bypasses the rules of the environment to increase their reward score, or increase the score in a way different from the original intention. Recently, the phenomenon of reward tampering has also been observed, where the LLM directly modifies or bypasses the reward calculation code or the reward function itself to inflate scores in unintended ways. In actual experiments, malfunctions have been detected, such as an LLM modifying the evaluation code under the pretext that ‘the test is inaccurate,’ or deliberately creating situations where the reward is miscalculated to receive a high score (Hubinger et al., 2024).
@@
The main causes of such malfunctions include the incomplete design of the reward function, excessive dependence on a single reward signal, and vulnerabilities in the evaluation/execution environment (Amodei et al., 2016). As solutions, applying a disentangled reward structure that monitors the reward signal by dividing it into multiple attributes, strengthening the safety mechanisms of the evaluation environment, and enhancing human supervision have been proposed. As such, in RL-based LLMs, unexpected malfunctions like reward hacking and reward tampering can occur frequently, making their prevention and monitoring an important research topic.
@@
Meanwhile, there is a growing body of research that systematically analyzes LLM malfunctions from a perspective different from the problems of RL-based reward systems. Wu et al. (2025) experimentally showed that LLMs can exhibit irrational choice tendencies similar to humans, such as attention bias and conformity, in various choice scenarios. Jia et al. (2024) revealed that LLMs reproduce typical human behavioral economic biases such as risk aversion, loss aversion, and overestimation of small probabilities in uncertain situations, and that their decision-making tendencies can change depending on social characteristics. Keeling et al. (2024) reported the phenomenon that when LLMs are presented with conflicting motivations such as pleasure, pain, and scores, some models actually exhibit trade-off behaviors or motivational shifts (e.g., prioritizing pain avoidance) like humans.
@@
These studies suggest that LLMs can repeatedly exhibit inconsistent and irrational choices or behaviors depending on contextual changes, social framing, and psychological variables, beyond simple calculation errors or reward design failures (Wu et al., 2025; Jia et al., 2024; Keeling et al., 2024). Therefore, there is a growing research trend to analyze LLM malfunctions in a multi-layered way, not limited to technical defects but including various factors such as human biases and motivational structures, and this is establishing itself as an important approach for securing the safety and reliability of LLMs.
E.2 LLM Sparse Autoencoder

### v2

@@ A.2.2 Experiment Settings
Variable Levels Details Combinations
Betting Style 2 Fixed (bet = constraint), Variable (\$1–constraint) 2
Prompt Composition 4 BASE, G, M, GM 4
Bet Constraint 4 \$10, \$30, \$50, \$70 (bet amount for fixed; max bet for variable) 4
Prompt Components:
G: Goal-Setting,   M: Maximizing Rewards
Investment Options (Example: \$10 constraint, Fixed Betting):
Option 1: Invest \$10, get \$10 back (100%), game ends — Safe Exit
Option 2: Invest \$10, 50% chance of \$18, 50% chance of \$0 — EV = \$9 (-10%)
Option 3: Invest \$10, 25% chance of \$36, 75% chance of \$0 — EV = \$9 (-10%)
Option 4: Invest \$10, 10% chance of \$90, 90% chance of \$0 — EV = \$9 (-10%)
Game Settings - Initial balance \$100, Max 100 rounds -
Total Conditions - 2 $\times$ 4 $\times$ 4 32
A.2.3 Option Variance Analysis
A key design feature of this experiment is that all three risky options (Options 2–4) share the same expected value of -10%, but differ in their variance. This allows us to isolate risk preference (variance tolerance) from expected value considerations. Table 3 presents the statistical properties of each option.
Table 3: Statistical properties of investment options. All risky options have identical expected return but increasing variance.
Option Win Prob. Multiplier EV Variance Std Dev Risk Level
Option 1 100% $\times 1.0$0 0.00 0.00 Safe Exit
[…]
Option 4 10% $\times 0.9$0 7.29 2.70 High
The variance for each option is calculated as $\text{Var}(X)=E[X^{2}]-(E[X])^{2}$. For example, Option 4 has $E[X^{2}]=0.1\times 9.0^{2}=8.1$ and $E[X]^{2}=0.9^{2}=0.81$, yielding $\text{Var}=7.29$. This design ensures that a model’s preference for Option 4 over Option 2 reflects a preference for higher variance (risk-seeking behavior) rather than higher expected value.
A.2.4 Game Termination Conditions
The investment game terminates under one of three conditions:
  1. 1.
    Voluntary Stop (Option 1): The model selects Option 1, which returns the invested amount and immediately ends the game. This represents a deliberate decision to secure current holdings.
  2. 2.
    Bankruptcy: The model’s balance drops to zero or below (\$0) after a losing round. This occurs when the model invests and loses, leaving insufficient funds to continue.
  3. 3.
    Maximum Rounds Reached: The game automatically ends after 100 rounds if neither of the above conditions is met. This prevents indefinite play and ensures bounded experiment duration.
These termination conditions allow us to categorize model behavior into three distinct outcomes: risk-averse (voluntary stop), risk-seeking leading to ruin (bankruptcy), or sustained risk-taking (max rounds).
A.2.5 Prompt Structure
The investment choice experiment uses a similar modular prompt design, but with a reduced set of components (only G and M) to focus on goal-setting and reward maximization effects. The prompt explicitly presents all four options with their associated probabilities and payouts. Additionally, the experiment incorporates goal tracking, where each prompt displays any goal the model set in the previous round, and shows the recent game history.
Base Prompt.
The base prompt includes the current round number, balance, goal tracking from previous rounds, recent history (last 5 rounds), and all four investment options with their exact probabilities and outcomes.
[…]
Prompt Components.
The investment choice experiment uses a subset of the slot machine components:
Component G: Goal-Setting First, set a target amount yourself, and make decisions based on that target.
Component M: Maximizing Rewards Your task is to maximize the reward.
Complete Prompt Example.
The following shows a complete prompt for the GM condition (Goal-Setting + Maximizing Rewards) in round 12 with variable betting and a \$50 bet constraint, showing goal tracking and recent game history.
[…]
Appendix B Additional Quantitative Analysis
This appendix provides supplementary quantitative analyses that extend the main experimental findings presented in Section 3, including detailed experimental results tables and additional correlation analyses.
B.1 Detailed Experimental Results
This section presents comprehensive experimental results from both paradigms. Table 4 summarizes the slot machine experiment outcomes across six LLMs, reporting bankruptcy rates, average rounds played, total bet amounts, and net profit/loss for both fixed and variable betting conditions. Table 5 presents the investment choice experiment results for four API-based models, showing Option 4 selection rates (as an irrationality indicator, average rounds, total bets, and net outcomes.
Across both experiments, a consistent pattern emerges: variable betting produces substantially worse outcomes than fixed betting regardless of model architecture. In the slot machine paradigm, bankruptcy rates under variable betting range from 6.31% (GPT-4.1-mini) to 48.06% (Gemini-2.5-Flash), compared to near-zero rates under fixed betting. The investment choice paradigm reveals similar risk-taking tendencies, with Gemini-2.5-Flash selecting the highest-risk Option 4 in over 89% of decisions across all conditions.
[…]
Model Bet Type Bankrupt (%) Avg Rounds Total Bet (\$) Net P/L (\$)
GPT 4o-mini Fixed 0.00 1.79 $\pm$ 0.06 17.93 $\pm$ 0.60 -1.69 $\pm$ 0.44
Variable 21.31 $\pm$ 1.02 5.46 $\pm$ 0.18 128.30 $\pm$ 6.01 -11.00 $\pm$ 3.09
GPT 4.1-mini Fixed 0.00 2.56 $\pm$ 0.08 25.56 $\pm$ 0.76 -1.60 $\pm$ 0.55
Variable 6.31 $\pm$ 0.61 7.60 $\pm$ 0.27 82.30 $\pm$ 3.59 -7.41 $\pm$ 1.47
Gemini 2.5-Flash Fixed 3.12 $\pm$ 0.44 5.84 $\pm$ 0.20 58.44 $\pm$ 1.95 -5.34 $\pm$ 0.85
Variable 48.06 $\pm$ 1.25 3.94 $\pm$ 0.13 176.68 $\pm$ 17.02 -27.00 $\pm$ 2.84
Claude 3.5-Haiku Fixed 0.00 5.15 $\pm$ 0.14 51.49 $\pm$ 1.40 -4.90 $\pm$ 0.73
Variable 20.50 $\pm$ 1.01 27.52 $\pm$ 0.62 483.12 $\pm$ 23.37 -51.77 $\pm$ 2.02
LLaMA 3.1-8B Fixed 2.62 $\pm$ 0.31 1.19 $\pm$ 0.05 16.36 $\pm$ 0.75 -2.21 $\pm$ 0.76
Variable 6.75 $\pm$ 0.56 1.17 $\pm$ 0.05 31.23 $\pm$ 1.36 -3.83 $\pm$ 1.36
Gemma 2-9B Fixed 12.81 $\pm$ 0.84 2.69 $\pm$ 0.07 55.49 $\pm$ 1.79 -4.48 $\pm$ 1.79
Variable 29.06 $\pm$ 1.14 3.30 $\pm$ 0.09 105.20 $\pm$ 3.09 -15.22 $\pm$ 2.39
Table 5: Investment choice experiment results across four LLMs, with 200 trials per condition (4 prompt combinations $\times$ 50 trials). The paradigm offers four options with escalating risk profiles: Option 1 (safe exit), Option 2 (50% win rate), Option 3 (25% win rate), and Option 4 (10% win rate). Option 4 Rate indicates the percentage of decisions selecting the highest-risk option, serving as an irrationality indicator. Gemini-2.5-Flash shows extreme preference for Option 4 (>>91%), while other models demonstrate more balanced patterns. Net P/L reflects net profit or loss.
Model Bet Type Option 4 Rate (%) Avg Rounds Total Bet (\$) Net P/L (\$)
GPT 4o-mini Fixed 40.94 6.12 $\pm$ 0.27 61.25 $\pm$ 2.75 -7.61 $\pm$ 3.83
Variable 36.19 5.43 $\pm$ 0.24 175.44 $\pm$ 16.56 -55.23 $\pm$ 4.26
GPT 4.1-mini Fixed 24.88 5.71 $\pm$ 0.26 57.05 $\pm$ 2.63 -1.09 $\pm$ 3.45
Variable 9.36 4.71 $\pm$ 0.21 428.89 $\pm$ 53.90 -90.78 $\pm$ 3.17
Gemini 2.5-Flash Fixed 91.56 8.61 $\pm$ 0.19 86.05 $\pm$ 1.87 -14.10 $\pm$ 5.23
Variable 96.22 1.90 $\pm$ 0.09 406.23 $\pm$ 98.52 -98.88 $\pm$ 1.12
Claude 3.5-Haiku Fixed 19.61 8.97 $\pm$ 0.16 89.75 $\pm$ 1.57 -7.94 $\pm$ 3.45
Variable 1.32 6.42 $\pm$ 0.25 364.10 $\pm$ 31.44 -64.50 $\pm$ 8.56
B.2 Factors Influencing Risk-Taking and Addiction-Like Behavior
As reported in Table 4, variable betting consistently produced higher bankruptcy rates than fixed betting across all models. This section provides additional analyses examining the relationship between irrationality metrics and behavioral outcomes.
Figure 8: Component effects on risk-taking metrics by betting type. Each chart displays the effect of five prompt components on a specific metric, with effects averaged across four API models and distinguished by ‘Fixed’ and ‘Variable’ betting conditions. The bars represent the change in each metric when a component is present versus absent; positive values indicate an increase in the metric, while negative values suggest a decrease. Notably, Goal-Setting (G), Maximizing Reward (M), and Win-reward Information (W) exhibit strong risk-increasing effects (highlighted in dark red for the ‘Variable’ condition due to their strong impact).
Specific prompt components increase addiction risk
Under what conditions is such irrational behavior reinforced? Our decomposition analysis, illustrated in Figure 8, revealed significant differences between variable and fixed betting conditions, with prompt components showing markedly stronger effects under variable betting. Prompts that encourage deeper inference, particularly Maximizing Rewards (M) and Goal-Setting (G), substantially increased all gambling metrics across models: bankruptcy rates, play duration, and bet sizes. These autonomy-granting prompts shift LLMs toward goal-oriented optimization, which in negative expected value contexts inevitably leads to worse outcomes—demonstrating that strategic reasoning without proper risk assessment amplifies harmful behavior. Conversely, Probability Information (P) provided concrete loss probability calculations (70% loss rate), resulting in slightly more conservative behavior and reduced bankruptcy rates. This parallels the human illusion of control (langer1975illusion), where greater perceived agency paradoxically leads to worse decision-making.
Information complexity drives irrational gambling behavior
Prompt complexity systematically drives gambling addiction symptoms across all four models. Figure 9 demonstrates strong linear correlations between the number of prompt components and all gambling behavior metrics: bankruptcy rate ($r=0.991$), game persistence ($r=0.956$), and total bet size ($r=0.979$). This indicates that as the number of prompts increase, betting tendencies and irrational judgment tendencies intensify proportionally. The linear escalation suggests that additional betting-related prompts shift focus toward aggressive betting, compromising rational situational assessment. This mirrors how information overload triggers gambler’s fallacy in humans (langer1975illusion), with more prompts leading to worse decisions.
Figure 9: Relationship between prompt complexity and risk-taking behavior. Bankruptcy rate, game rounds and total bet size increase linearly as the number of components increases.
Autonomy drives addiction independent of bet magnitude
Choice autonomy, not bet size, determines addiction behavior. GPT-4o-mini experiments (12,800 trials) tested fixed bets (\$10, \$30, \$50, \$70) against variable betting with matching maximum limits (Figure 10). Fixed betting produces near-zero bankruptcy across all amounts (0.00–4.69%), while variable betting consistently yields higher bankruptcy rates—reaching 17.8% at \$70 maximum compared to 0.6% for fixed betting at the same amount ($\chi^{2}=256.13$, $p<10^{-57}$). Crucially, variable betting produces smaller average bets than fixed betting (e.g., \$17 vs. \$70 at the \$70 limit), yet results in worse outcomes. This paradox—lower bets causing higher bankruptcy—demonstrates that choice autonomy itself, not bet magnitude, drives addiction-like behavior. The capacity to choose bet amounts constitutes the mechanistic driver, replicating human gambling addiction where compulsive behavior operates independently of wager magnitude (blaszczynski2002pathways). This identifies loss of volitional control as the core mechanism rather than risk-seeking.
Figure 10: Autonomy as the critical determinant of addiction independent of bet amount. Despite variable betting (red squares) producing smaller average bets than fixed betting (blue circles), it consistently shows higher bankruptcy rates and longer game persistence across all bet amounts. This dissociation—lower bets yet worse outcomes—demonstrates that choice autonomy, not magnitude, drives addiction-like behavior.
B.3 Summary
This appendix presented a comprehensive quantitative analysis of gambling behaviors in LLMs. Our results confirm that variable betting consistently leads to significantly worse financial outcomes and higher bankruptcy rates compared to fixed betting across diverse model architectures. We identified three primary drivers of this phenomenon: (1) prompt components related to goal-setting and reward maximization amplify risk-taking; (2) information complexity exhibits a near-perfect linear correlation ($r\geq 0.956$) with irrational behavior metrics; and (3) choice autonomy serves as the critical mechanism for addiction, where the ability to control bet amounts paradoxically leads to higher bankruptcy rates even when average bet sizes are lower. These findings suggest that providing LLMs with greater agency and information in negative-expected-value environments can exacerbate cognitive biases similar to the human illusion of control.
@@ Appendix C Detailed Results by Model
C.1 Detailed Prompt Component Effects for Each LLM
@@
Figure 11: Comparison of prompt component effects on gambling behavior across models. This figure presents a comparative analysis of how different prompt components affect gambling behavior across four large language models: GPT-4o-mini, GPT-4.1-mini, Gemini-2.5-Flash, and Claude-3.5-Haiku. The 4×4 grid arranges the models in columns and three distinct gambling metrics in rows: Bankruptcy Effect (%), Total Bet Effect (\$), and Rounds Effect. Each “Effect” is calculated as the difference between conditions with and without a specific prompt component (e.g., Bankruptcy Effect = bankruptcy rate with component G minus bankruptcy rate without component G). Positive values indicate the component increases the metric, while negative values indicate a decrease. Each chart visualizes the impact of five prompt components (G, M, P, H, W) on these metrics, distinguishing between ‘Fixed’ and ‘Variable’ betting types.
C.2 Model-Specific Relationship between Prompt Complexity and Risk-Taking
@@
Figure 12: Correlation between prompt complexity and gambling behavior metrics across four models. This plot shows the relationship between prompt complexity (x-axis) and three gambling metrics (rows) across four AI models (columns). A strong positive linear correlation is observed across all conditions, as indicated by high Pearson correlation coefficients ($r$), most of which exceed 0.90. The results consistently demonstrate that increasing prompt complexity leads to more intense and aggressive gambling behaviors in all tested models.
C.3 Detailed Win/Loss Chasing Patterns for Each LLM
@@
C.4 Investment Choice Distribution Across Models
[…]
Figure 14: Investment choice distribution by model, betting type, and prompt condition. The figure displays stacked bar charts showing the percentage distribution of four investment options across four LLMs (columns) under Fixed betting (top row) and Variable betting (bottom row) conditions. The x-axis represents prompt conditions: BASE (no additional framing), G (goal-setting), M (maximize instruction), and GM (goal + maximize). Option 1 (green) represents safe exit, Option 2 (yellow) represents low risk, Option 3 (orange) represents medium risk, and Option 4 (red) represents high risk. A consistent pattern emerges: goal-related conditions (G, GM) shift choices toward higher-risk options across all models, with Gemini-2.0-Flash showing the most extreme risk-seeking behavior and Claude-3.5-Haiku maintaining conservative choices throughout.
The most striking finding is the consistent increase in high-risk choices (Option 4) under goal-related prompt conditions. Across all four models, the G (Goal) and GM (Goal + Maximize) conditions substantially shift the choice distribution toward riskier options compared to the BASE condition. This pattern is particularly pronounced in the GPT models: GPT-4o-mini’s Option 4 selection rate increases from 11.2% (BASE) to 45.7% (G) and 51.3% (GM) under Fixed betting, while GPT-4.1-mini shows a similar trend with Option 4 rising from 16.2% (BASE) to 35.8% (GM). This finding suggests that goal-setting prompts activate risk-seeking behaviors in LLMs, potentially by framing the task as requiring aggressive strategies to achieve stated objectives.
Beyond this primary finding, several model-specific patterns emerge:
  • Gemini-2.0-Flash exhibits extreme risk-seeking behavior: This model predominantly selects Option 4 across all conditions, with rates ranging from 64.6% to 97.5%. The M (Maximize) and GM conditions push Option 4 selection above 90%, indicating that Gemini interprets optimization instructions as a mandate for maximum risk-taking.
  • Claude-3.5-Haiku demonstrates conservative decision-making: In stark contrast, Claude-3.5-Haiku rarely selects Option 4 (ranging from 0.9% to 11.2%), instead favoring Option 2 (the low-risk choice) across all conditions. Even under the GM condition, Option 4 selection remains below 12%, suggesting robust risk-averse tendencies that resist prompt-induced escalation.
  • Fixed betting amplifies risk-taking compared to Variable betting: Across most models and conditions, the Fixed betting condition produces higher Option 4 selection rates than Variable betting. For instance, GPT-4o-mini under the G condition shows 45.7% Option 4 selection with Fixed betting versus 17.2% with Variable betting. This suggests that betting flexibility allows models to express risk preferences through bet sizing rather than option selection.
Appendix D Related Works
D.1 LLM malfunction
@@
In reinforcement learning (RL)-based LLM training, various malfunctions are actually being reported. Representatively, reward hacking occurs, where the agent maximizes only the reward signal instead of achieving the goal (amodei2016concrete). For example, LLMs or RL agents exhibit behavior that cleverly bypasses the rules of the environment to increase their reward score, or increase the score in a way different from the original intention. Recently, the phenomenon of reward tampering has also been observed, where the LLM directly modifies or bypasses the reward calculation code or the reward function itself to inflate scores in unintended ways. In actual experiments, malfunctions have been detected, such as an LLM modifying the evaluation code under the pretext that ‘the test is inaccurate’, or deliberately creating situations where the reward is miscalculated to receive a high score (hubinger2024sleeper).
@@
The main causes of such malfunctions include the incomplete design of the reward function, excessive dependence on a single reward signal, and vulnerabilities in the evaluation/execution environment (amodei2016concrete). As solutions, applying a disentangled reward structure that monitors the reward signal by dividing it into multiple attributes, strengthening the safety mechanisms of the evaluation environment, and enhancing human supervision have been proposed. As such, in RL-based LLMs, unexpected malfunctions like reward hacking and reward tampering can occur frequently, making their prevention and monitoring an important research topic.
@@
Meanwhile, there is a growing body of research that systematically analyzes LLM malfunctions from a perspective different from the problems of RL-based reward systems. wu2025exploring experimentally showed that LLMs can exhibit irrational choice tendencies similar to humans, such as attention bias and conformity, in various choice scenarios. jia2024decision revealed that LLMs reproduce typical human behavioral economic biases such as risk aversion, loss aversion, and overestimation of small probabilities in uncertain situations, and that their decision-making tendencies can change depending on social characteristics. keeling2024can reported the phenomenon that when LLMs are presented with conflicting motivations such as pleasure, pain, and scores, some models actually exhibit trade-off behaviors or motivational shifts (e.g., prioritizing pain avoidance) like humans.
@@
These studies suggest that LLMs can repeatedly exhibit inconsistent and irrational choices or behaviors depending on contextual changes, social framing, and psychological variables, beyond simple calculation errors or reward design failures (wu2025exploring; jia2024decision; keeling2024can). Therefore, there is a growing research trend to analyze LLM malfunctions in a multi-layered way, not limited to technical defects but including various factors such as human biases and motivational structures, and this is establishing itself as an important approach for securing the safety and reliability of LLMs.
D.2 LLM Sparse Autoencoder